tigerlover (tigerlover):

Suppose a triangle has sides a, b, and c, and that a2 + b2 < c2. Let be the measure of the angle opposite the side of length c. Which of the following must be true? Check all that apply. A. is an acute angle. B. The triangle is a right triangle. C. cos < 0 D. The triangle is not a right triangle.

OpenStudy (mathstudent55):

Some info is missing. What is the measure of the angle opposite the side of length c? In A. What is an acute angle? In C. cos of what?

tigerlover (tigerlover):

hold on

tigerlover (tigerlover):

\[con \theta < 0\]

tigerlover (tigerlover):

\[\theta is an \acute \angle\]

OpenStudy (mathstudent55):

I'm sure you know the Pythagorean theorem and its famous equation: \(a^2 + b^2 = c^2\)

OpenStudy (mathstudent55):

Here you have \(a^2 + b^2 < c^2\). Is this triangle a right triangle?

tigerlover (tigerlover):

it doesnt give a picture

tigerlover (tigerlover):

its just the question and answer

OpenStudy (mathstudent55):

The Pythagorean theorem states that if you have a right triangle with legs a and b, and hypotenuse c, then the equation below is true. \(a^2 + b^2 = c^2\) Also, the converse of the Pythagorean theorem is true. In other words, if you have a triangle with side lengths a, b, and c, and the equation \(a^2 + b^2 = c^2 \) is true, then the triangle is a right triangle.

OpenStudy (mathstudent55):

You are given the inequality \(a^2 + b^2 < c^2\) Can this triangle be a right triangle?

tigerlover (tigerlover):

yes?

OpenStudy (mathstudent55):

The problem clearly states that for this triangle a^2 + b^2 < c^2 In a right triangle a^2 + b^2 = c^2 must be true. Can a^2 + b^2 = c^2 and a^2 + b^2 < c^2 be both true at the same time?

tigerlover (tigerlover):

i figuered it out.

OpenStudy (mathstudent55):

Which choices did you choose?